OpenStudy (anonymous):

Find the inverse of the function. f(x)=^2+4/3

OpenStudy (ipwnbunnies):

I assume it's f(x) = x^2 + 4/3

OpenStudy (ipwnbunnies):

To the find the inverse, let's rewrite the equation, replacing f(x) with y. y = x^2 + 4/3

OpenStudy (ipwnbunnies):

Then, swap the places of x and y. So you'll get x = ... Then, solve for y again. Finally, replace y with f^(-1)(x), which means inverse of a function.

OpenStudy (ipwnbunnies):

Oh, ok. f(x) = (x^2 + 4)/3 Same deal, can you follow my steps?

OpenStudy (anonymous):

okay let me see

OpenStudy (anonymous):

okay so you said swap the places of x and y..so 3/x^2+4?

OpenStudy (anonymous):

i mean o x=3/x^2+4

OpenStudy (ipwnbunnies):

Err, no try again. So we have this: \[y = \frac{x^2 + 4}{3}\]

OpenStudy (ipwnbunnies):

This is what I mean by swap places: \[x = \frac{y^{2} + 4}{3}\]

OpenStudy (anonymous):

ohh okay so umm whats next?

OpenStudy (ipwnbunnies):

Solve for y. We need to get this in the correct form for the inverse notation.

OpenStudy (anonymous):

i forgot how to solve.. do i just subtract 3 with 4 and y^2?

OpenStudy (ipwnbunnies):

No. Let's get rid of the 3 in the denominator first. How can you do that?

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

do i switch it again?

OpenStudy (ipwnbunnies):

No, we're not switching it again. We're just solving for y. How about multiplying both sides by 3?

OpenStudy (anonymous):

okay so i multiply 4*3 which is 12 and what about the y^2?

OpenStudy (ipwnbunnies):

That is incorrect. Do you not know how to simplify fractions? \[x(3) = (\frac{y^2 + 4}{3})(3)\] \[3x = y^2 + 4\]

OpenStudy (anonymous):

no i dont. i really suck at this. okay so you subtract 3 on each side or no?

OpenStudy (anonymous):

i mean for the next step you did

OpenStudy (ipwnbunnies):

I multiplied both sides by 3. It will cancel on the right side because 3/3 = 1. And x will just turn into 3x. Now, on the newest equation, how can we get y by itself?

OpenStudy (ipwnbunnies):

Is there something you can subtract?

OpenStudy (anonymous):

is it 4?

OpenStudy (ipwnbunnies):

Yes, rewrite the equation now.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

wait do i subtract 4 on the 3x?

OpenStudy (anonymous):

3x=y^2+4 -4 -4

OpenStudy (anonymous):

like that?

OpenStudy (ipwnbunnies):

Yes yes. What do you get now.

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

i mean 1 lol

OpenStudy (anonymous):

1=y^2

OpenStudy (ipwnbunnies):

No, you can't subtract terms with different variables. 3x can only be added or subtracted by terms with x in them.

OpenStudy (ipwnbunnies):

\[3x - 4 = y^2\] Do you know how to get rid of the power on the y now?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh wait hold on

OpenStudy (anonymous):

so is y^2 subtracted by 4 then? and 4 and 4 are crossed out

OpenStudy (ipwnbunnies):

Yeah, on the right side. The 4's got cancelled out because we subtracted 4 from both sides.

OpenStudy (ipwnbunnies):

Do you know about square roots?

OpenStudy (anonymous):

yes

OpenStudy (ipwnbunnies):

Good. Square root both sides so we can get rid of the squared term on the y.

OpenStudy (anonymous):

square y^2? and 3?

OpenStudy (anonymous):

im sorry but you could show me how you did it

OpenStudy (anonymous):

nvm idk..

OpenStudy (ipwnbunnies):